Mathematics 244 – Linear Algebra Definitions Quiz 9 – April 20, 2007

Your Name:

A) (5) Define: B is a basis for the vector space V.

B) (5) Say whether the following statement is true or false and justify your claim: If A is a diagonalizable n × n matrix, then all of its eigenvalues are nonzero.